An object of mass 24kg is accelerated up a frictionless plane inclined at an angle of 37⁰ with the horizontal by a constant force .starting at the bottom from the rest, it covers a distance of 18m in 3 seconds
1) what is the average power required to accomplish the process ?
2)what is the instantaneous power required at the end of the 3 second interval ?
1
Expert's answer
2021-08-23T12:40:46-0400

Find the resultant acceleration:


"a=\frac{2d}{t^2}."


Find the net force:


"F=ma=\frac{2md}{t^2}."

Find the average velocity:


"v=\frac dt."


1) The average power is the force times average speed:


"P=\frac{2md^2}{t^3}=576\text { W}."

2) Find the final speed at the end of 3rd second:


"v_f=at=\frac{2d}{t}."


The instantaneous power is


"P_i=Fv_f=\frac{4md^2}{t^3}=2P=1152\text{ W}."